OpenStudy (anonymous):

3(-2 + 4)^3 - 2 (-4 + 1^2

OpenStudy (anonymous):

bracket after 1

OpenStudy (anonymous):

3(2)^3-2(-3)=30

OpenStudy (anonymous):

3*8 - 2*18 = -12

OpenStudy (anonymous):

could you show your work ?

OpenStudy (anonymous):

is last part -2(-4+1)^2 ?

OpenStudy (anonymous):

-4+1^2=-3

OpenStudy (anonymous):

it's - 2(-4 + 1)^2

OpenStudy (anonymous):

it's not negative it's minus

OpenStudy (anonymous):

right wel thats - 2 * (-3)^2 = -2 * 9 = -18

OpenStudy (anonymous):

could you show how you got the answer, im kind of confused :S

OpenStudy (anonymous):

first bit is 3 * 2^3 = 3 * 8 = 24 so 24 - 18 = 6 sorry - i was wrong first time

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank -you !

OpenStudy (anonymous):

yw